If 20ml of 0.1M NaOH is added to 30ml of .0.2M ch3cooh apka=4.74,the pH of a resulting solution
1)4.44
2)9.56
3)8.96
4)9.26

Grade:12th pass

1 Answers

Arun
5 years ago

Number of gram equivalents(eq) = moles x acidity of a base (or bascity of an acid.)

Or

= Molarity x Volume in litres x acidity of a base (or bascity of an acid)

So,

Eq(NaOH) = 20/1000 x 0.1 x 1 = 0.002 eq

Eq(CH3COOH) = 30/1000 x0.2 x1 = 0.006 eq

That means acid = base. Hence a salt would be formed in the solution with total volume = 30 + 20 = 50 ml

The pH of a strong base weak acid salt solution =

7 + 0.5(pKa +logC) where C is concentration of salt

concentration of salt = eq/volume =1000 x 0.005/50

= 0.01

So

pH = 7 + 0.5(4.74 +log0.01)

pH = 7 + 1.37 = 8.37